• Calculate the concentration of OH - ions in a 1.4 x 10-3 M HCl solution.
  • Calculate the pH of the buffer system made up of 0.15 M NH3 and 0.35 M NH4Cl.
1
Expert's answer
2022-05-13T07:53:04-0400
  • Calculate the concentration of OH - ions in a 1.4 x 10-3 M HCl solution.

pOH = 14 - pH

pH = - log [HCl] = - log 1.4 x 10-3  =2.85

pOH = 14 -2.85 = 11.15

[OH-] = 7.08 х 10 -12


  • Calculate the pH of the buffer system made up of 0.15 M NH3 and 0.35 M NH4Cl.

pOH= pKb + log[salt]/[Acid]

 Kb of NH3 is 1.8x10-5

pKb= -logKb = -log(1.8x10-5) = 4.745

pOH= 4.745 + log(0.35/0.15) = 4.38

pH = 14 - pOH = 14 - 4.38 = 9.62